OpenStudy (anonymous):

suppose a bike and rider weigh 100kg. bike is accelerated at rest to 20km/h in 10 seconds, what is the acceleration? show the force required is approximately 55.5N

OpenStudy (callisto):

20km/h = 20x1000/(60x60) = 5.5556 ms^-1 acceleration = (v-u)/t = (5.5556-0)/10 = 0.55556ms^-2 By Newton's 2nd law ΣF =ma = 100x0.55556 = 55.6N ( difference is due to the way you round the number) Got it?

OpenStudy (anonymous):

ah right, i didn't remember that a=(v-u)/t thanks yeah, i got that ") now on to the next section:)